# Cooperative Task Spaces for Multi-Arm Manipulation Control based on Similarity Transformations

### v1

@@ I Introduction
Humans excel at dexterous manipulation using the redundancy of their arms and hands to achieve complex tasks where cooperative behaviors between the different kinematic chains is required. Many environments are built around these advanced manipulation capabilities, and robots that are increasingly operating within them. Yet, it still remains challenging for robots to achieve dexterous manipulation that is on par with humans [19]. Nevertheless, robots are now expected to handle various tasks that require them to interact with objects in a way that goes beyond traditional parallel jaw grippers or suction cups that are often featured on single arm manipulators. These types of end-effectors are limiting the robot manipulation capabilities, since many objects are often more complex and require different strategies. For instance, big and bulky objects may call for dual-arm manipulators for picking and placing. This kind of coordination between two arms is vital when grasping, lifting, and transferring objects of varying shape, weight, and fragility. This is similarly true for intricate tasks such as in-hand manipulating and multi-fingered grasping using robotics hands. The common ground here is coordinating the movements of two or more parallel kinematic chains to interact with the environment during manipulation tasks.
@@
However, many manipulation tasks do not require exploiting the full redundancy of the system. This is because the kinematic chains are not acting independently but instead are coordinated to achieve a shared objective. As a result, the effective behavior of the system often lies on a lower-dimensional manifold embedded within the full configuration space. We refer to these manifolds exhibiting specific geometric structures as cooperative task spaces. Recognizing and leveraging this structure not only reduces the dimensionality of the control or planning problems but also improves interpretability. Moreover, these geometric structures naturally give rise to geometric nullspaces, which can be exploited for secondary objectives, such as regulating contact forces, without interfering with the primary task objective, as we previously demonstrated in [4].
@@
Existing work using dual quaternions [1] and geometric algebra [23] has begun to explore cooperative task spaces, particularly in the context of dual-arm manipulation. However, the notion of cooperation extends well beyond dual-arm settings. Many real-world scenarios involving multiple kinematic chains, such as torso-arm-hand coordination or multi-fingered in-hand manipulation, can benefit from a unified geometric framework that generalizes cooperative control. To illustrate this, we provide a non-exhaustive collection of robotic systems and manipulation challenges that can be abstracted through cooperative subspaces corresponding to geometric primitives, as shown in Figure 1. The geometric primitives only depend on the number of kinematic chains that are involved in the task. Hence, they model the cooperative behavior, and are not chosen based on an approximation of the object shape. Our overarching goal is to generalize operational space control [20], which was originally formulated for single manipulators, to systems with multiple kinematic chains by representing cooperative task spaces as geometric primitives within the control architecture.
@@
In this article, we address the challenge of cooperative manipulation control for robotic systems involving multiple parallel kinematic chains. We do so by combining the various ideas around cooperative control, grasping nullspaces and geometric primitives, into a single, mathematically coherent framework by introducing cooperative geometric primitives and integrating them into geometrically consistent control schemes. For this, we leverage conformal geometric algebra (CGA) to find algebraic objects that represent geometric primitives that go beyond keypoints. In prior work, we have shown how these primitives define geometric nullspaces and how they can be used in optimal control [22]. In the same optimal control framework, we have then extended the cooperative dual-task space using CGA for bimanual robots [23]. Here, we further extend this idea, by exploring how cooperative geometric primitives enable decentralized control strategies that harmonize task execution, disturbance rejection, and inter-agent coordination. We show that by exploiting CGA, different systems can be modelled in a uniform manner that allows informed collaboration strategies based on the geometric primitives and conformal transformations that can be expressed in this algebra. Our contributions therefore are
@@ I-A Related Work
A common example for highlighting the challenges of cooperative manipulation is the collective transport of big and bulky objects that cannot be handled by a single robot. This task is very challenging, since multiple robots interact with the same object via local control actions [8], which requires them to cooperate in order to achieve a common, global goal. Traditional approaches often use a leader-follower strategy, where the different roles are predefined, and the agents communicate explicitly. However, these strategies often fail due to uncertainties stemming from the task parameters, environmental dynamics, or the team composition [9]. Solving these issues requires strategies that do not rely on centralized coordination, but instead use the manipulated object as a physical channel for implicit coordination. By sensing the exerted forces on the shared payload, the robots can decompose the motion into an allowed task space motion and a nullspace motion that is used rejecting perturbations [6]. This strategy minimizes force conflicts, and enables cooperative manipulation even under strict communication constraints. Recent advances in cooperative control favor frameworks that further emphasize distributed optimization and adaptability. They optimize the individual performance of the robots in the system that suffers reduced operation capabilities due to the physical connection through grasping [17]. Furthermore, the optimization can include objectives for enhancing the manipulability, avoiding obstacles, and reducing internal forces caused by robot motion errors due to coupling [16]. Hence, there is an inherent trade-off between physical coupling (e.g. rigid grasps) and operational flexibility, which makes it important to deal with disturbances to avoid unexpected behaviors [2]. Learning-based approaches try to simplify the problem of coordinating multiple agents by introducing actions through shared latent spaces, which reduces the sample complexity of high-dimensional tasks like dual-arm manipulation [3]. Alternatively, bi-manual end-effector poses are learned from demonstrations using task-parameterized dynamical systems [33].
@@
Dynamic Movement Primitives (DMPs) have emerged as a powerful way to encode complex trajectories while preserving stability guarantees. A comprehensive tutorial survey frames DMPs as a bridge between model‑based control and data‑driven learning, highlighting their suitability for multi‑robot settings where each agent can generate locally consistent motions that nevertheless respect a globally defined attractor [28]. Building on this foundation, it has been demonstrated that DMPs can be combined with velocity scaling to personalise human‑robot collaborative transport. By modulating the DMP speed according to the measured interaction forces, the system achieves smooth, safe hand‑overs even when the human operator varies the payload weight or trajectory on‑the‑fly [11]. Other work has leveraged human motion prediction based on DMPs to enable seamless collaborative object transfer, showing that anticipating the human partner’s trajectory improves coordination and reduces interaction forces [30]. The sensed forces on the shared object become the scaling signal for the DMPs, thereby turning the payload itself into a communication medium that implicitly synchronises the robots’ motion. While many cooperative manipulation studies assume rigid payloads, real‑world logistics increasingly involve deformable or soft objects (e.g., bags of grain, flexible containers). Movement‑Primitive Diffusion introduced a generative framework that learns gentle manipulation policies for deformable items directly from demonstrations [29]. The diffusion model produces smooth DMP parameters that respect material compliance, effectively reducing internal stresses that would otherwise arise from hard‑grasping strategies. This approach complements the earlier discussion on internal‑force minimisation [16] by providing a data‑driven means to shape the primitive itself, rather than relying solely on post‑hoc optimisation. Moreover, the diffusion‑generated primitives can be shared across robots, enabling distributed execution without explicit communication—each robot samples the same stochastic primitive conditioned on its local perception of the object’s deformation.
@@
Some approaches are inspired by the formation control that is often used in the area of autonomous ground vehicles [34], and are extended to manipulation, where the agents are treated as part of a cohesive geometric formation [31]. In this setting, one of the agents can be a human that is guiding the robots, which in turn try to maintain desired geometry w.r.t. each other, while reducing internal stresses [32]. It can then be shown that, the internal interactions forces in cooperative manipulation that arise from a formation of robots, lie in the null space of the grasp matrix, i.e. the dynamics of the object are not affected by them [7]. This insight also bridges the connection to literature on grasping and in-hand manipulation, where the problem setting is similar, in the sense that multiple parallel kinematic chains are tasks to cooperatively manipulate a shared object. Both scenarios have a high kinematic redundancy can be exploited [35]. We can therefore also mirror control strategies for robotic grip adjustment, to enable adaptive force modulation for cooperative robotic manipulation in response to environmental feedback. Here, compliant control behaviors are achieved through impedance control schemes on the object level [27]. Dexterity in contact-rich manipulation is often limited, because in contrast to humans, typically grasping models for robotic hands only use single contact points at the fingertips, instead of the whole surface [5]. Smart gripper design can alleviate this issue to a certain extent, where the geometry is optimized to achieve more dexterous manipulation capabilities, such as fixing the movement of the object’s pose along sphere [25]. This highlights the importance of geometric considerations for encoding spatial relationships and constraints and the role of geometric primitives.
@@
Geometric primitives also play an important role in recent advances on general manipulation control, currently mainly in the form of object-centric keypoints. The goal here is to reduce complex tasks across a category of objects, where shape variations are allowed, to geometric constraints on keypoints [14]. When used in closed-loop manipulation control for achieving contact-rich tasks, this keypoint-based representation leads to automatic generalization to a category of objects by incorporating an object-centric action representation [24]. The simplicity of the representation via keypoints has also been proven to have advantages for visual imitation learning. Here, task representations are extracted from demonstrations and decomposed into keypoint-based geometric constraints [13]. This approach has been extended to bimanual manipulation and was shown to generalize to novel scenes in-category objects, due to the object-centric, embodiment-independent, and viewpoint-invariant representation [12]. Expressing task constraints via geometric constraints on keypoints also facilitates the connection to vision-language models, such that tasks can be defined using language instructions [18].
@@ II-A1 Geometric Primitives
This means that the set of conformalized Euclidean points that result in zero using the outer product forms the geometric primitives [26]. We have shown previously how this outer product formulation can be formulated within an optimal control framework to define reaching tasks involving the geometric primitives [22].
@@ II-B Robot Modeling using CGA
where $B_{j}$ is the bivector describing essentially the screw axis of the $j$-th joint. We described the derivation of the analytic $\bm{\mathcal{J}}^{A}_{\bm{q}}$ and geometric $\bm{\mathcal{J}}^{G}_{\bm{q}}$ Jacobians in [22]. Expressed w.r.t. the end-effector motor, their relationship can be found as
@@ II-C Geometric Algebra for Optimal Control
In [22], we presented an approach for optimal control using CGA that used the outer product for distance computations between geometric primitives. For a conformal point $P(\bm{q})$, this distance can be formulated as the outer product of the target geometric primitive and the point
@@ II-D Geometric Nullspace for Impedance Control
Nullspaces offer an intuitive way to formulate hierarchical control objectives, i.e. to introduce secondary objectives that do not disturb the tracking of the primary one. We have previously exploited the geometric nullspace formed by a line primitive to separate the objectives of tracking a line, while exerting a force along that line in a tactile control application [4]. This separation was achieved purely geometrically without the need to derive an appropriate precision matrix and transforming it to the correct reference frame. Hence, the geometric nullspaces are entirely coordinate-free and thus greatly simplify the definition of hierarchical control objectives.
@@ II-E Cooperative Dual Task Space
In this section, we briefly introduce the cooperative dual-task space using CGA that we presented in [23] as an extension of [1]. The formulation of the CDTS uses two motors to express the collaborative behaviour of the two manipulators. The motors represent a relative and an absolute pose. We first find the relative motor as
@@
where $\bm{q}_{1}$ and $\bm{q}_{2}$ denote the joint configurations of the first and second manipulator, respectively. $M_{1}$ and $M_{2}$ are the corresponding end-effector motors found using the forward kinematics formula from Equation (7). For brevity, we omit the derivation of the corresponding analytic Jacobians $\bm{\mathcal{J}}^{A}_{r}(\bm{q}_{1},\bm{q}_{2})$ and $\bm{\mathcal{J}}^{A}_{a}(\bm{q}_{1},\bm{q}_{2})$ and instead refer readers to [23] for details.
@@ III-C Similarity Transformations
Its bivector $B_{S}(X_{1},X_{2})$ found from the logarithmic map can then be used in the optimization problem given in Equation (39). A general formula to find a versor $V_{C}(X_{1},X_{2})$ that transforms $X_{1}$ to $X_{2}$ was derived in [21]
@@
where $c$ is a normalization constant that depends on the geometric primitives. However, using this general formula, $V_{C}(X_{1},X_{2})\in\mathcal{C}$, i.e. the resulting versor will be an element of the conformal transformation group $\mathcal{C}$, not of the similarity transformation group $\mathcal{S}$. In order to restrict the versor to $\mathcal{S}$, we present the necessary derivations that give a similarity transformation between pairs of geometric primitives according to Equation (20). Note that, this clearly shows that the transformation between primitives is not unique, as was already stated in [21].
@@ III-J Geometric Nullspace
In our work, we use the cooperative similarity transformations to extend the notion of end-effector poses to systems of multiple kinematic chains. The geometric nullspace is induced by the cooperative geometric primitive, since by definition of the geometric primitives, they form a nullspace under the outer product as seen from Equation (3). In our formulation of a cooperative geometric primitive, we use the end-effector points of the involved kinematic chains as its defining points. Consequently, those end‑effector points can move arbitrarily along directions tangent to the cooperative primitive without altering the primitive itself. Hence, the cooperative geometric primitive directly encodes the system’s geometric nullspace. Due to geometric algebra, the primitives are algebraic objects and not simply parameterizations. But the explicit parameterization, such as radius and center of a sphere, can easily be obtained from these algebraic objects. Furthermore, we can easily relate the derivatives of those parameters to the joint angles by using the analytic Jacobian of the cooperative geometric primitive $\bm{\mathcal{J}}_{c}^{A}(\bm{q})$ that we showed in Equation (15). This enables using recent redundancy resolution methods such as [15] or [10].

### v2

@@ I Introduction
Humans excel at dexterous manipulation using the redundancy of their arms and hands to achieve complex tasks where cooperative behaviors between the different kinematic chains is required. Many environments are built around these advanced manipulation capabilities, and robots that are increasingly operating within them. Yet, it still remains challenging for robots to achieve dexterous manipulation that is on par with humans [undefr]. Nevertheless, robots are now expected to handle various tasks that require them to interact with objects in a way that goes beyond traditional parallel jaw grippers or suction cups that are often featured on single arm manipulators. These types of end-effectors are limiting the robot manipulation capabilities, since many objects are often more complex and require different strategies. For instance, big and bulky objects may call for dual-arm manipulators for picking and placing. This kind of coordination between two arms is vital when grasping, lifting, and transferring objects of varying shape, weight, and fragility. This is similarly true for intricate tasks such as in-hand manipulating and multi-fingered grasping using robotics hands. The common ground here is coordinating the movements of two or more parallel kinematic chains to interact with the environment during manipulation tasks.
@@
However, many manipulation tasks do not require exploiting the full redundancy of the system. This is because the kinematic chains are not acting independently but instead are coordinated to achieve a shared objective. As a result, the effective behavior of the system often lies on a lower-dimensional manifold embedded within the full configuration space. We refer to these manifolds exhibiting specific geometric structures as cooperative task spaces. Recognizing and leveraging this structure not only reduces the dimensionality of the control or planning problems but also improves interpretability. Moreover, these geometric structures naturally give rise to geometric nullspaces, which can be exploited for secondary objectives, such as regulating contact forces, without interfering with the primary task objective, as we previously demonstrated in [undefc].
@@
Existing work using dual quaternions [undef] and geometric algebra [undefv] has begun to explore cooperative task spaces, particularly in the context of dual-arm manipulation. However, the notion of cooperation extends well beyond dual-arm settings. Many real-world scenarios involving multiple kinematic chains, such as torso-arm-hand coordination or multi-fingered in-hand manipulation, can benefit from a unified geometric framework that generalizes cooperative control. To illustrate this, we provide a non-exhaustive collection of robotic systems and manipulation challenges that can be abstracted through cooperative subspaces corresponding to geometric primitives, as shown in Figure 1. The geometric primitives only depend on the number of kinematic chains that are involved in the task. Hence, they model the cooperative behavior, and are not chosen based on an approximation of the object shape. Our overarching goal is to generalize operational space control [undefs], which was originally formulated for single manipulators, to systems with multiple kinematic chains by representing cooperative task spaces as geometric primitives within the control architecture.
@@
In this article, we address the challenge of cooperative manipulation control for robotic systems involving multiple parallel kinematic chains. We do so by combining the various ideas around cooperative control, grasping nullspaces and geometric primitives, into a single, mathematically coherent framework by introducing cooperative geometric primitives and integrating them into geometrically consistent control schemes. For this, we leverage conformal geometric algebra (CGA) to find algebraic objects that represent geometric primitives that go beyond keypoints. In prior work, we have shown how these primitives define geometric nullspaces and how they can be used in optimal control [undefu]. In the same optimal control framework, we have then extended the cooperative dual-task space using CGA for bimanual robots [undefv]. Here, we further extend this idea, by exploring how cooperative geometric primitives enable decentralized control strategies that harmonize task execution, disturbance rejection, and inter-agent coordination. We show that by exploiting CGA, different systems can be modelled in a uniform manner that allows informed collaboration strategies based on the geometric primitives and conformal transformations that can be expressed in this algebra. Our contributions therefore are
@@ I-A Related Work
A common example for highlighting the challenges of cooperative manipulation is the collective transport of big and bulky objects that cannot be handled by a single robot. This task is very challenging, since multiple robots interact with the same object via local control actions [undefg], which requires them to cooperate in order to achieve a common, global goal. Traditional approaches often use a leader-follower strategy, where the different roles are predefined, and the agents communicate explicitly. However, these strategies often fail due to uncertainties stemming from the task parameters, environmental dynamics, or the team composition [undefh]. Solving these issues requires strategies that do not rely on centralized coordination, but instead use the manipulated object as a physical channel for implicit coordination. By sensing the exerted forces on the shared payload, the robots can decompose the motion into an allowed task space motion and a nullspace motion that is used rejecting perturbations [undefe]. This strategy minimizes force conflicts, and enables cooperative manipulation even under strict communication constraints. Recent advances in cooperative control favor frameworks that further emphasize distributed optimization and adaptability. They optimize the individual performance of the robots in the system that suffers reduced operation capabilities due to the physical connection through grasping [undefp]. Furthermore, the optimization can include objectives for enhancing the manipulability, avoiding obstacles, and reducing internal forces caused by robot motion errors due to coupling [undefo]. Hence, there is an inherent trade-off between physical coupling (e.g. rigid grasps) and operational flexibility, which makes it important to deal with disturbances to avoid unexpected behaviors [undefa]. Learning-based approaches try to simplify the problem of coordinating multiple agents by introducing actions through shared latent spaces, which reduces the sample complexity of high-dimensional tasks like dual-arm manipulation [undefb]. Alternatively, bi-manual end-effector poses are learned from demonstrations using task-parameterized dynamical systems [undefaf].
@@
Dynamic Movement Primitives (DMPs) have emerged as a powerful way to encode complex trajectories while preserving stability guarantees. A comprehensive tutorial survey frames DMPs as a bridge between model‑based control and data‑driven learning, highlighting their suitability for multi‑robot settings where each agent can generate locally consistent motions that nevertheless respect a globally defined attractor [undefaa]. Building on this foundation, it has been demonstrated that DMPs can be combined with velocity scaling to personalise human‑robot collaborative transport. By modulating the DMP speed according to the measured interaction forces, the system achieves smooth, safe hand‑overs even when the human operator varies the payload weight or trajectory on‑the‑fly [undefj]. Other work has leveraged human motion prediction based on DMPs to enable seamless collaborative object transfer, showing that anticipating the human partner’s trajectory improves coordination and reduces interaction forces [undefac]. The sensed forces on the shared object become the scaling signal for the DMPs, thereby turning the payload itself into a communication medium that implicitly synchronises the robots’ motion. While many cooperative manipulation studies assume rigid payloads, real‑world logistics increasingly involve deformable or soft objects (e.g., bags of grain, flexible containers). Movement‑Primitive Diffusion introduced a generative framework that learns gentle manipulation policies for deformable items directly from demonstrations [undefab]. The diffusion model produces smooth DMP parameters that respect material compliance, effectively reducing internal stresses that would otherwise arise from hard‑grasping strategies. This approach complements the earlier discussion on internal‑force minimisation [undefo] by providing a data‑driven means to shape the primitive itself, rather than relying solely on post‑hoc optimisation. Moreover, the diffusion‑generated primitives can be shared across robots, enabling distributed execution without explicit communication—each robot samples the same stochastic primitive conditioned on its local perception of the object’s deformation.
@@
Some approaches are inspired by the formation control that is often used in the area of autonomous ground vehicles [undefag], and are extended to manipulation, where the agents are treated as part of a cohesive geometric formation [undefad]. In this setting, one of the agents can be a human that is guiding the robots, which in turn try to maintain desired geometry w.r.t. each other, while reducing internal stresses [undefae]. It can then be shown that, the internal interactions forces in cooperative manipulation that arise from a formation of robots, lie in the null space of the grasp matrix, i.e. the dynamics of the object are not affected by them [undeff]. This insight also bridges the connection to literature on grasping and in-hand manipulation, where the problem setting is similar, in the sense that multiple parallel kinematic chains are tasks to cooperatively manipulate a shared object. Both scenarios have a high kinematic redundancy can be exploited [undefah]. We can therefore also mirror control strategies for robotic grip adjustment, to enable adaptive force modulation for cooperative robotic manipulation in response to environmental feedback. Here, compliant control behaviors are achieved through impedance control schemes on the object level [undefz]. Dexterity in contact-rich manipulation is often limited, because in contrast to humans, typically grasping models for robotic hands only use single contact points at the fingertips, instead of the whole surface [undefd]. Smart gripper design can alleviate this issue to a certain extent, where the geometry is optimized to achieve more dexterous manipulation capabilities, such as fixing the movement of the object’s pose along sphere [undefx]. This highlights the importance of geometric considerations for encoding spatial relationships and constraints and the role of geometric primitives.
@@
Geometric primitives also play an important role in recent advances on general manipulation control, currently mainly in the form of object-centric keypoints. The goal here is to reduce complex tasks across a category of objects, where shape variations are allowed, to geometric constraints on keypoints [undefm]. When used in closed-loop manipulation control for achieving contact-rich tasks, this keypoint-based representation leads to automatic generalization to a category of objects by incorporating an object-centric action representation [undefw]. The simplicity of the representation via keypoints has also been proven to have advantages for visual imitation learning. Here, task representations are extracted from demonstrations and decomposed into keypoint-based geometric constraints [undefl]. This approach has been extended to bimanual manipulation and was shown to generalize to novel scenes in-category objects, due to the object-centric, embodiment-independent, and viewpoint-invariant representation [undefk]. Expressing task constraints via geometric constraints on keypoints also facilitates the connection to vision-language models, such that tasks can be defined using language instructions [undefq].
@@ II-A1 Geometric Primitives
This means that the set of conformalized Euclidean points that result in zero using the outer product forms the geometric primitives [undefy]. We have shown previously how this outer product formulation can be formulated within an optimal control framework to define reaching tasks involving the geometric primitives [undefu].
@@ II-B Robot Modeling using CGA
where $B_{j}$ is the bivector describing essentially the screw axis of the $j$-th joint. We described the derivation of the analytic $\bm{\mathcal{J}}^{A}_{\bm{q}}$ and geometric $\bm{\mathcal{J}}^{G}_{\bm{q}}$ Jacobians in [undefu]. Expressed w.r.t. the end-effector motor, their relationship can be found as
@@ II-C Geometric Algebra for Optimal Control
In [undefu], we presented an approach for optimal control using CGA that used the outer product for distance computations between geometric primitives. For a conformal point $P(\bm{q})$, this distance can be formulated as the outer product of the target geometric primitive and the point
@@ II-D Geometric Nullspace for Impedance Control
Nullspaces offer an intuitive way to formulate hierarchical control objectives, i.e. to introduce secondary objectives that do not disturb the tracking of the primary one. We have previously exploited the geometric nullspace formed by a line primitive to separate the objectives of tracking a line, while exerting a force along that line in a tactile control application [undefc]. This separation was achieved purely geometrically without the need to derive an appropriate precision matrix and transforming it to the correct reference frame. Hence, the geometric nullspaces are entirely coordinate-free and thus greatly simplify the definition of hierarchical control objectives.
@@ II-E Cooperative Dual Task Space
In this section, we briefly introduce the cooperative dual-task space using CGA that we presented in [undefv] as an extension of [undef]. The formulation of the CDTS uses two motors to express the collaborative behaviour of the two manipulators. The motors represent a relative and an absolute pose. We first find the relative motor as
@@
where $\bm{q}_{1}$ and $\bm{q}_{2}$ denote the joint configurations of the first and second manipulator, respectively. $M_{1}$ and $M_{2}$ are the corresponding end-effector motors found using the forward kinematics formula from Equation (7). For brevity, we omit the derivation of the corresponding analytic Jacobians $\bm{\mathcal{J}}^{A}_{r}(\bm{q}_{1},\bm{q}_{2})$ and $\bm{\mathcal{J}}^{A}_{a}(\bm{q}_{1},\bm{q}_{2})$ and instead refer readers to [undefv] for details.
@@ III-C Similarity Transformations
Its bivector $B_{S}(X_{1},X_{2})$ found from the logarithmic map can then be used in the optimization problem given in Equation (39). A general formula to find a versor $V_{C}(X_{1},X_{2})$ that transforms $X_{1}$ to $X_{2}$ was derived in [undeft]
@@
where $c$ is a normalization constant that depends on the geometric primitives. However, using this general formula, $V_{C}(X_{1},X_{2})\in\mathcal{C}$, i.e. the resulting versor will be an element of the conformal transformation group $\mathcal{C}$, not of the similarity transformation group $\mathcal{S}$. In order to restrict the versor to $\mathcal{S}$, we present the necessary derivations that give a similarity transformation between pairs of geometric primitives according to Equation (20). Note that, this clearly shows that the transformation between primitives is not unique, as was already stated in [undeft].
@@ III-J Geometric Nullspace
In our work, we use the cooperative similarity transformations to extend the notion of end-effector poses to systems of multiple kinematic chains. The geometric nullspace is induced by the cooperative geometric primitive, since by definition of the geometric primitives, they form a nullspace under the outer product as seen from Equation (3). In our formulation of a cooperative geometric primitive, we use the end-effector points of the involved kinematic chains as its defining points. Consequently, those end‑effector points can move arbitrarily along directions tangent to the cooperative primitive without altering the primitive itself. Hence, the cooperative geometric primitive directly encodes the system’s geometric nullspace. Due to geometric algebra, the primitives are algebraic objects and not simply parameterizations. But the explicit parameterization, such as radius and center of a sphere, can easily be obtained from these algebraic objects. Furthermore, we can easily relate the derivatives of those parameters to the joint angles by using the analytic Jacobian of the cooperative geometric primitive $\bm{\mathcal{J}}_{c}^{A}(\bm{q})$ that we showed in Equation (15). This enables using recent redundancy resolution methods such as [undefn] or [undefi].